OpenStudy (anonymous):

Solve the following system of equations. 2x + 3y – z = 1 3x + y + 2z = 12 x + 2y – 3z = –5

sam (.sam.):

You can do simultaneous equations few times or use vector, x=3,y=-1,z=2

OpenStudy (anonymous):

Okay, thank you

sam (.sam.):

2x + 3y – z = 1 x = (1- 3y + z )/2..................(1) ---------------------- 3 (1- 3y + z )/2 + y + 2z = 12 (3(1-3y+z))/(2)+y+2z=12 (3(-3y+z+1))/(2)+y+2z=12 (3(-3y+z+1))/(2)+y*(2)/(2)+2z*(2)/(2)=12 (3(-3y+z+1))/(2)+(2y)/(2)+(4z)/(2)=12 (3(-3y+z+1)+(2y)+(4z))/(2)=12 (-7y+7z+3)/(2)=12 -7y+7z+3=24 -7y=-7z+21 y=z-3.....................(2) ------------------------------- From eqn 3, (1- 3(z-3) + z )/2 + 2(z-3) – 3z = –5 (1-3(z-3)+z)/(2)+2(z-3)-3z=-5 (1-3z+9+z)/(2)+2(z-3)-3z=-5 (10-2z)/(2)+2(z-3)-3z=-5 (2(5-z))/(2)+2(z-3)-3z=-5 (5-z)+2(z-3)-3z=-5 5-z+2(z-3)-3z=-5 5-z+2z-6-3z=-5 -1-2z=-5 -2z=-4 z=2...................(3) ======================== Sub (3) to (2) y=-1 Sub (3) and (2) to (1) x=3 ======================== x=3,y=-1,z=2